OpenStudy (neo92):

how to estimate the number of iterations needed in bisection method in order to have the answer correct to given number of decimal places?

OpenStudy (anonymous):

hi do you have books on this subject

OpenStudy (neo92):

u mean ebuks?

OpenStudy (anonymous):

no, school books

OpenStudy (neo92):

yep. but it dosent xplain much abt da above ques. bt i gotta do it 4 ny assignment

OpenStudy (anonymous):

are you in algebra or in numerical analysis?

OpenStudy (neo92):

numerical analysis

OpenStudy (anonymous):

ah ok,,,,,

OpenStudy (neo92):

can u hrlp me????

OpenStudy (anonymous):

[Bisection Method Bisection Method Algorithm: Input \epsilon (i.e level of accuracy) Set a, b, (The interval) Compute midpoint=\frac{a+b}{2} If f(midpoint)=0, midpoint=root If f(midpoint)>0 (i.e function is positive at the midpoint), replace midpoint with the endpoint where function was positive If f(midpoint)<0 (i.e negative), replace midpoint with endpoint where function was negative If (b-a)>\epsilon (if root is not found and level of desired accuracy is not reached, repeat 3-6)

OpenStudy (anonymous):

The equation t3 + 4 t2 - 1 = 0 has a positive root r in [0,1]. f(0)<0 and f(1)>0. Since f(0.5) = 0.125 the root is in [0, 0.5]. Since f(0.25) = -0.73 the root is in [0.25, 0.5]. Since f(0.375) = -0.38 the root is in [0.375, 0.5]. Since f(0.4375) = -0.15 the root is in [0.4375, 0.5]. Since f(0.46875) = -0.018 the root is in [0.46875, 0.5]. Since f(0.484375) = 0.05 the root is in [0.46875, 0.484375]. ... and so we approach the root 0.472834

OpenStudy (neo92):

how can i knw da number of iterations needed b4 solving it?

OpenStudy (anonymous):

take a good guess first

OpenStudy (anonymous):

if f(c) is positive on your first guess, then on your 2nd guess changes sign then the root is in between

OpenStudy (anonymous):

do you have a sample problem there?

OpenStudy (neo92):

yep wait i'l post it here

OpenStudy (anonymous):

use your calculator,,ok post them

OpenStudy (neo92):

It is required to find \[\sqrt[3]{2}\] to five decimal places, by solving the eqn \[x^3-2=0\] by means of the method of bisection. Estimate the number of iterations that will be required, and show this is so by performing the iterations.

OpenStudy (anonymous):

f(2)=6 f(1)==1 here you know its between 1 and 2, so (1+2)/2=1.5 now try 1.5

OpenStudy (anonymous):

f(1.5)=1.375 so (1+1.5)/2=1.25

OpenStudy (neo92):

yeah i get it. but it says 'estimate the number of iterations needed' ??????

OpenStudy (anonymous):

yes count your number of iteration till you get closer to the root

OpenStudy (neo92):

:( didnt get it?

OpenStudy (anonymous):

The method selects the subinterval that is a bracket as a new interval to be used in the next step. In this way the interval that contains a zero of f is reduced in width by 50% at each step. The process is continued until the interval is sufficiently small.

OpenStudy (anonymous):

The method is guaranteed to converge to a root of f if f is a continuous function on the interval [a, b] and f(a) and f(b) have opposite signs. The absolute error is halved at each step so the method converges linearly, which is comparatively slow. Specifically, if p1 = (a+b)/2 is the midpoint of the initial interval, and pn is the midpoint of the interval in the nth step, then the difference between pn and a solution p is bounded by

OpenStudy (anonymous):

absolute value of[Pn-P]< or =abs[b-a]/2^n

OpenStudy (anonymous):

That formula can be used to determine in advance the number of iterations that the bisection method would need to converge to a root to within a certain tolerance.

OpenStudy (neo92):

thnks a lot..

OpenStudy (anonymous):

is it clearer now?

OpenStudy (neo92):

yep.. i was wandering how to find da number of iteratins b4 solvng.. my txt buk dosent mention abt it

OpenStudy (anonymous):

Initialization: The bisection method is initialized by specifying the function f(x), the interval [a,b], and the tolerance > 0. We also check whether f(a) = 0 or f(b) = 0, and if so return the value of a or b and exit. Loop: Let m = (a + b)/2 be the midpoint of the interval [a,b]. Compute the signs of f(a), f(m), and f(b). If any are zero, return the corresponding point and exit. Assuming none are zero, if f(a) and f(m) have opposite sides, replace b by m, else replace a by m. If the length of the [a,b] is less than , return the value of a and exit. Analysis: When we enter the loop f(a) and f(b) have opposite sign. It follows that either f(m) and f(a) have opposite sign or f(m) and f(b) have opposite sign. Thus the initial conditions are still satisfied each time we enter the loop. The length of the initial interval is (b - a). After one time through the loop the length is (b - a)/2, after two times it is (b - a)/4, and after n passes through the loop, the length of the remaining interval is (b - a)/2n. No matter how small , eventually (b - a)/2^n <E . In fact we can solve this inequality for n: (b - a)/2^n <E 2^n >(b - a)/E n ln 2 >ln(b - a) - ln(E) n>[ln(b - a) - ln(E)]/ln 2. Thus the algorithm terminates after at most M passes through the loop where M is the first integer larger than [ln(b - a) - ln(E)]/ln 2.

OpenStudy (anonymous):

Starting with the interval [1,2], find sqrt(2) to within two decimal places (to within an error of .01).

OpenStudy (anonymous):

try that for yourself for practice

OpenStudy (anonymous):

solve f(x) = x^2 -2.

OpenStudy (anonymous):

the final result is the approximation 1.41406 for the sqrt(2).

OpenStudy (neo92):

wait il do it n post da answer here

OpenStudy (anonymous):

make a table that have like this: a b b-a m=(a+b)/2 f(a) f(b) f(m)

OpenStudy (anonymous):

answer it like this: a b b-a m=(a+b)/2 f(a) f(b) f(m) 1) 1 2 1 1.5 -1 2 25 2) 3) 5) 6) 7) 8)1.41406 1.42187 .0078125 -.0004343 .0217285

OpenStudy (neo92):

bro i gt 2 go nw.. can gve me ur email so i cn mail it 2 u? or shall i post it here when im back?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

or you can post em if you want

OpenStudy (neo92):

ok..thnkx

OpenStudy (anonymous):

yw good luck now